The weights (in kg) of 11 Stem B follow a normal distribution and has a mean of 50 and a standard deviation of 5. How many students have weights greater than 57?

1
Expert's answer
2022-03-10T13:31:11-0500

"P(X>57)=P(Z>\frac{57-50}{5})=P(Z>1.4)=1-P(Z<1.4)=0.0808."

"N=11*0.0808\approx 1."